OpenStudy (anonymous):

Which value is equivalent to 8 multiplied by 4 multiplied by 2 whole over 8 multiplied by 7, the whole raised to the power of 2 multiplied by 8 to the power of 0 over 7 to the power of negative 3, whole to the power of 3 multiplied by 7 to the power of negative 9? 64 over 49 8 over 49 16 over 7 512 over 7

OpenStudy (anonymous):

@AihberKhan

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

@THEHELPER123

OpenStudy (anonymous):

@harleygirl18

OpenStudy (anonymous):

@surjithayer

OpenStudy (aihberkhan):

Okay let's do this step by step.

OpenStudy (aihberkhan):

\(8 * 4 * 2 = 64\)

OpenStudy (aihberkhan):

\(64 / (8*7) = 56\)

OpenStudy (aihberkhan):

\(\large56^2 = \frac{ 64 }{ 49 }\)

OpenStudy (aihberkhan):

That's it! :)

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) A medal would also be appreciated! Just click best response next to my name! Also, a fan would be appreciated as well! Just hover over my icon and click "Become A Fan". This will allow you to see every time I am online! :) If you see that I am online and need help with a question, just tag me in your question! @justen472